OpenStudy (anonymous):

There are 6 girls and 7 boys in a class. A team of 10 players is to be selected from the class. If the selection is random, what is the probability of selecting a team of 5 girls and 5 boys?

OpenStudy (anonymous):

This may help you:

OpenStudy (anonymous):

\[P=\frac{ c_{5}^{6}*_c{5}^{7} }{ c _{10}^{13} }=?\]

OpenStudy (anonymous):

\[P=\frac{ c_{5}^{6}*c _{5}^{7} }{ c _{10}^{13} }\]

OpenStudy (anonymous):

remember\[c _{r}^{n}=c _{n-r}^{n}\]

OpenStudy (anonymous):

\[c _{5}^{6}=c _{6-5}^{6}=c _{1}^{6}=\frac{ 6 }{ 1 }=6\] \[c _{5}^{7}=c _{7-5}^{7}=c _{2}^{7}=\frac{ 7*6 }{ 2*1 }=21\] \[c _{10}^{13}=c _{13-10}^{13}=c _{3}^{13}=\frac{ 13*12*11 }{ 3*2*1 }=286\] find p

OpenStudy (anonymous):

I have problems comprehending equations like that. Is there a simpler way to do it?